# Distributed Online Feedback Optimization for Real-time Distribution System Voltage Regulation

### v3

@@ Abstract
We investigate the real-time voltage regulation problem in distribution systems employing online feedback optimization (OFO) with short-range communication between physical neighbours. OFO does not need an accurate grid model nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties and disturbances, which render it particularly suitable for real-time distribution system applications. However, many OFO controllers require centralized communication, making them susceptible to single-point failures. This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence. Numerical study results demonstrate that the proposed design achieves effective voltage regulation and outperforms other distributed and local approaches.
@@ I Introduction
OFO also brings distinct advantages over other existing approaches. Compared to local droop control [11, 3, 12], OFO ensures more reliable voltage regulation [3] and can pursue grid-level objectives. In contrast to deep reinforcement learning [13, 14], OFO does not need a complicated (often centralized offline) training process while offering theoretical guarantees to obtain globally optimal solutions. Furthermore, unlike using traditional voltage regulators such as on-load tap changers and capacitor banks [15], OFO provides rapid voltage regulation with inverter-interfaced DERs and can complement these slow-reacting devices in real time [16].
@@
While many OFO controllers rely on centralized communication [1, 17, 2, 8, 4, 5, 18, 9, 19, 20, 21], distributed OFO based on communication between physical neighbours appears to be a compelling approach due to its robustness to single-point failures[6, 3, 16, 22, 23, 24]. Among the existing distributed designs, [6, 3, 16] demand particular forms of objective functions, while [22] relies on disseminating global information through neighbouring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In [23], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes. This heuristic approach does not assure adequate voltage regulation.
@@
In this paper, we propose a distributed OFO design inspired by [24], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e. $\mathbf{X}^{-1}$ is sparse. Both our proposed and their approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint. It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information. However, as shown in [25], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus convergence. With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [25]. Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design. As a remedy, [24] introduced additional dualization of the DER capacity constraint on top of the Euclidean projection to ensure the DER constraint satisfaction. Nevertheless, this strategy takes a prohibitively large number of iterations to converge, which may jeopardize its performance in online implementation.
@@
We tackle this challenge by proposing an iterative algorithm to solve this non-Euclidean projection problem that requires only local information and operates in an online feedback-based manner. This leads to our overall nested feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, while the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints using local information. To summarize, the main contributions of this paper are:
  • A local iterative algorithm is proposed to solve the non-Euclidean projection problem that operates in an online feedback-based manner.
  • An efficient nested feedback optimization strategy that ensures adequate voltage regulation while eliminating the need for an accurate grid model and consumption of non-controllable loads is developed.
  • The proposed strategy is theoretically analyzed and numerically assessed on a 96-bus system with high-resolution generation and load data.
@@ II-A Distribution system modeling
$\mathbf{R}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}r_{hk},\mathbf{X}_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}x_{hk}.$ (3)
@@
$\mathbf{R}_{ij}^{-1}\neq 0,\mathbf{X}_{ij}^{-1}\neq 0\iff(i,j)\in\mathcal{E}\text{ or }i=j.$ (5)
@@ II-C Centralized primal-dual gradient projection algorithm
Solving (6) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be readily available. The feedforward approach also lacks robustness [7]. Thus, this section introduces a feedback controller based on OFO. The controller leverages primal-dual gradient projection (PDGP) as the underlying algorithm [1, 17, 2] to track the time-varying optimizers of (6), which requires a centralized gather-and-broadcast communication architecture. Specifically, the PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
@@ III-B Proposed distributed approach
To tackle this challenge, this paper proposes an iterative algorithm to solve (13) that requires only local information and is implemented in an online feedback-based manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (13) to generate feasible setpoints. Our key observation is that given a feasible $\mathbf{u}^{\tau}\in\mathcal{X}^{k}$, the gradient of the objective function in (13) at $\mathbf{u}^{\tau}$ is $\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})$, which can be approximated by the voltage difference $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ according to (2) assuming system disturbances do not change significantly when all nodes are locally implementing this inner iterative algorithm, i.e.
$\mathbf{X}(\mathbf{u}^{\tau}-\mathbf{\dot{q}}^{k+1})\approx\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1}).$ (14)
The proposed algorithm is based on the non-scaled Euclidean gradient projection and proceeds as:
  1. 1.
    Implement
    $\mathbf{q}^{k}_{\epsilon}=\mathbf{q}^{k}+\epsilon(\mathbf{\dot{q}}^{k+1}-\mathbf{q}^{k}),$ (15)
    and measure the voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$. Estimate
    $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})\approx\mathbf{\tilde{v}}^{k}+(\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})-\mathbf{\tilde{v}}^{k})/\epsilon.$ (16)
  2. 2.
    Given a feasible $\mathbf{u}^{\tau}$, implement it, measure the voltage $\mathbf{\tilde{v}}(\mathbf{u}^{\tau})$, compute the next implementable reactive power setpoint $\mathbf{u}^{\tau+1}$ as (17), and repeat this step for $T$ iterations.
    $\mathbf{u}^{\tau+1}=[\mathbf{u}^{\tau}-\alpha^{u}(\mathbf{\tilde{v}}(\mathbf{u}^{\tau})-\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1}))]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}.$ (17)
Remark 1.
In the first step, $\epsilon$ represents a small exploration parameter, which is similar to that in model-free feedback optimization [29, 30, 31, 32]. Since $\mathbf{\dot{q}}^{k+1}$ can be pretty far away from the feasible set, it might be impossible to implement it. Therefore, we implement $\mathbf{q}^{k}_{\epsilon}$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$ based on it. This is illustrated in Fig. 4. This could still result in minor temporary DER capacity limit violations, which should be manageable by the DERs. To ensure strict capacity limit satisfaction, the original feasible set can be slightly deflated [32]. The initial feasible $\mathbf{u}^{\tau}$ in the second step can be chosen as $\mathbf{q}^{k}$. The online implementation of the overall nested algorithm is summarized in Algorithm 1.
Data: $\mathbf{q}^{0}$
Result: $\mathbf{q}^{k}$, $k\geq 0$; $\mathbf{u}^{\tau}$, $\tau\leq T$
1 Set k = 0;
@@ 3       Implement $\mathbf{q}^{k}$ and collect voltage measurement $\tilde{\mathbf{v}}(\mathbf{q}^{k})$;
5       Implement $\mathbf{q}^{k}_{\epsilon}$, measure voltage $\mathbf{\tilde{v}}(\mathbf{q}^{k}_{\epsilon})$ and estimate $\mathbf{\tilde{v}}(\mathbf{\dot{q}}^{k+1})$;
@@ 6       Set $\tau=0$;
8             Implement $\mathbf{u}^{\tau}$ and measure voltage $\tilde{\mathbf{v}}(\mathbf{u}^{\tau})$ (choose $\mathbf{u}^{0}$ as $\mathbf{q}^{k}$);
9             Calculate next implementable $\mathbf{u}^{\tau+1}$ via (17);
@@ III-C Convergence analysis
Based on the first assumption, $\mathbf{v}(\mathbf{u}^{\tau})-\mathbf{v}(\mathbf{\dot{q}}^{k+1})$ is the exact gradient of the objective function in (13) at $\mathbf{u}^{\tau}$. The inner loop then represents a standard Euclidean gradient projection algorithm to solve the quadratic program (13). Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [25] with $0<\alpha^{u}<2/\lambda_{max}(\mathbf{X})$ where $\lambda_{max}(\mathbf{X})$ is the largest eigenvalue of $\mathbf{X}$.
@@
The overall nested algorithm thus represents a scaled PDGP algorithm to solve (6). By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\sqrt{\frac{\alpha^{d}}{\alpha}}$ is the squared root of the ratio between the dual and primal step sizes, the problem (6) can be cast as an equivalent quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ as in (18).
@@
which are equivalent to (9a)-(9b) and (11)-(13) [25, Eqs. (2.37) and (2.38)]. Invoking Theorem 3 in [2] with $\alpha$ chosen using [2, Eq. (34)], (19) then converges to the optimizer of the regularized saddle-point problem of (18). Consequently, our proposed nested algorithm also converges to the optimizer of the regularized saddle-point problem of (6).
@@
First, while the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the actual voltages. Such an assumption was also made in the proofs of [3, 23]. Second, a large $T$ is not necessary in the simulation. We use $T=4$ so the inner loop might not converge. However, the algorithm performs well in the simulation. Notably, a similar assumption was made in [3].

### v4

# Link to IEEEXplore: https://ieeexplore.ieee.org/document/11066276

@@ Abstract
We investigate the real-time voltage regulation problem in distribution systems employing online feedback optimization (OFO) with short-range communication between physical neighbours. OFO does not need an accurate grid model nor estimated consumption of non-controllable loads, affords fast calculations, and demonstrates robustness to uncertainties and disturbances, which render it particularly suitable for real-time distribution system applications. However, many OFO controllers require centralized communication, making them susceptible to single-point failures. This paper proposes a distributed OFO design based on a nested feedback optimization strategy and analyzes its convergence. The strategy preserves end-users’ privacy by keeping voltage data local. Numerical study results demonstrate that the proposed design achieves effective voltage regulation and outperforms other distributed and local approaches.
@@ I Introduction
OFO also brings distinct advantages over other existing approaches. Compared to local droop control [11, 3, 12], OFO ensures more reliable voltage regulation [3] and can pursue grid-level objectives. In contrast to deep reinforcement learning [13, 14], OFO does not need a complicated (often centralized offline) training process while offering theoretical guarantees to obtain globally optimal solutions. Recent studies in [15, 16, 17, 18] have explored data-driven approaches to train linear and nonlinear local voltage controllers via deep learning, reinforcement learning, and unsupervised learning. However, these approaches demand extensive training data and lack optimality guarantees without real-time coordination between DERs. Furthermore, unlike using traditional voltage regulators such as on-load tap changers and capacitor banks [19], OFO provides rapid voltage regulation with inverter-interfaced DERs and can complement these slow-reacting devices in real time [20].
@@
While many OFO controllers rely on centralized communication [1, 21, 2, 8, 4, 5, 22, 9, 23, 24, 25], distributed OFO [6, 3, 20, 26, 27, 28] based on communication between physical neighbours appears to be a compelling approach due to its robustness to single-point failures and enhancement of privacy. Among the existing distributed designs, [6, 3, 20] demand particular forms of objective functions, while [26] relies on disseminating global information through neighbouring nodes. However, the systematic delays inherent in this information propagation may raise stability issues, potentially undermining its efficacy in practical distribution systems. In [27], the distributed design was achieved by only keeping components in the network sensitivity matrix related to local and neighbouring nodes. This heuristic approach does not assure adequate voltage regulation.
The motivation of this study is to develop a real-time voltage regulation strategy that optimally coordinates DERs on a fast timescale and operates without requiring an accurate grid model or load data. Furthermore, the strategy should be robust to single-point failures, meaning that unlike centralized approaches—where the failure of a single coordination unit can disrupt the entire system—the approach should ensure that voltage regulation remains functional even if some nodes fail. Finally, the strategy should preserve privacy. Exposing nodal voltage data could allow external entities to infer sensitive load information, which is typically considered private.
@@
In this paper, we propose a distributed OFO design that meets these requirements. The strategy is inspired by [28], where the distributed implementation was enabled by the sparsity of the inverse of the network sensitivity matrix $\mathbf{X}$, i.e. $\mathbf{X}^{-1}$ is sparse. Both our proposed and their approaches continue with scaling the gradient by $\mathbf{X}^{-1}$ followed by a projection step to ensure satisfaction of the DER capacity constraint. It is conceptually attractive to use the Euclidean distance in the projection, which can be implemented using only local information. However, as shown in [29], the distance measured by the norm $\lVert\mathbf{z}\rVert_{\mathbf{X}}=\sqrt{\mathbf{z}^{\intercal}\mathbf{X}\mathbf{z}}$ should be used instead in the projection step to ensure descent iterations and thus convergence. With the Euclidean distance, it would be a two-metric approach which can lead to algorithm divergence [29]. Solving this non-Euclidean projection problem would however require global information since $\mathbf{X}$ is dense, contradicting the distributed OFO design. As a remedy, [28] introduced additional dualization of the DER capacity constraint on top of the Euclidean projection to ensure the DER constraint satisfaction. Nevertheless, this strategy takes a prohibitively large number of iterations to converge, which may jeopardize its performance in online implementation.
@@
We tackle this challenge by proposing an iterative projected gradient descent algorithm to solve this non-Euclidean projection problem where each node only requires information from its neighbours. This leads to our overall nested feedback optimization approach, where the outer loop comprises the OFO iterations requiring communication between physical neighbours and yielding tentative but not necessarily feasible DER setpoints, while the inner loop solves the non-Euclidean projection problem to map the tentative setpoints to actual feasible setpoints. Unlike the centralized approaches, our method is robust to single-point failures as it operates without a central coordination unit and preserves privacy. Compared to [3, 6, 20], our approach accommodates a broader range of objective functions. In contrast to [26, 27], our approach keeps voltage information local enhancing privacy. Finally, compared to [28], our approach converges in significantly fewer iterations and achieves superior voltage regulation performance. To summarize, the main contributions of this paper are:
  • We develop a projected gradient descent algorithm that exchanges information between neighboring nodes in a backward-forward manner to solve the non-Euclidean projection problem.
  • We propose a nested feedback optimization strategy for distribution system voltage regulation that operates without requiring an accurate grid model or load data. The strategy optimally coordinates DERs on a fast timescale while ensuring robustness to single-point failures and preserving privacy by keeping voltage data local.
  • We theoretically prove the convergence of the proposed strategy. Simulation results demonstrate that it achieves comparable voltage regulation to its centralized counterpart while significantly outperforming other distributed and local approaches.
@@ II-A Distribution system modeling
$R_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}r_{hk},X_{ij}=\sum_{(h,k)\in\mathcal{E}_{i}\cap\mathcal{E}_{j}}x_{hk}.$ (3)
@@
$R_{ij}^{-1}\neq 0,X_{ij}^{-1}\neq 0\iff(i,j)\in\mathcal{E}\text{or }i=j.$ (5)
@@ II-B Problem formulation
Remark 1.
The objective function is designed to enable our distributed voltage regulation algorithm. Since the algorithm relies solely on local information and data from neighboring nodes, the gradient of the objective function must be locally computable. This requirement excludes certain objectives, such as loss minimization, which would require a centralized communication structure. Additionally, in case of overvoltage due to PV generation, minimizing reactive power usage contributes to loss reduction. When PV inverters absorb reactive power to lower voltage magnitudes, they increase reactive power transfer within the system, which already experiences reactive power consumption from non-controllable loads and cable losses. The more reactive power PV inverters absorb, the higher the system losses. Thus, minimizing their reactive power consumption helps reduce overall losses, as demonstrated by numerical results in Section IV-C. Finally, this objective function also enhances inverter reliability. As noted in [33], operating inverters at power factors significantly different from unity can reduce their lifespan. Therefore, our approach aligns with the interest of end-users by promoting longer inverter lifetime.
@@ II-C Centralized primal-dual gradient projection algorithm
Solving (6) offline in a feedforward way demands an accurate grid model and disturbance data to evaluate $v_{i}$, which might not be readily available. The feedforward approach also lacks robustness [7]. Thus, this section introduces a feedback controller based on OFO. The controller leverages primal-dual gradient projection (PDGP) as the underlying algorithm [1, 21, 2] to track the time-varying optimizer of (6), which requires a centralized gather-and-broadcast communication architecture. Compared to other dual optimization approaches such as dual ascent and the alternating method of multipliers (ADMM), PDGP offers additional tuning of primal step sizes, which is beneficial under fluctuating grid conditions, and enables faster implementation of DER setpoints. The centralized OFO-based solution is included primarily as a baseline for comparison and as a foundation for the distributed methods developed later.
Specifically, the PDGP algorithm performs projected gradient ascent and descent steps iteratively for the dual and primal variables, respectively. At each time step (iteration) $k$, it repeats the following three steps:
@@
Remark 2.
The regularization ensures strong convexity of the problem and guarantees convergence. However, it introduces a deviation from the exact solution of (7), resulting in a close approximation instead. As shown in [34], the discrepancy is bounded and proportional to $\sqrt{r_{d}}$. A practical guide for choosing the regularization parameters is to start with small values and gradually increase them until the algorithm stabilizes.
@@ III-B Proposed distributed approach
To tackle this challenge, this paper proposes an iterative projected gradient descent (PGD) algorithm to solve (13) that exchanges information only between physical neighbours in a backward-forward manner. This leads to our nested feedback optimization approach, where the outer loop comprises the OFO iterations giving tentative reactive power setpoints, and the inner loop solves the non-Euclidean projection problem (13) to generate feasible setpoints.
Given a $\mathbf{u}^{k,\tau}\in\mathcal{X}^{k}$ where $\tau=0,1,2,\cdots,T$ is the index for the inner loop, the PGD algorithm proceeds to compute $\mathbf{u}^{k,\tau+1}\in\mathcal{X}^{k}$ using (14) where $\mathbf{X}(\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1})$ is the gradient of the objective function of (13) at $\mathbf{u}^{k,\tau}$.
$\mathbf{u}^{k,\tau+1}=[\mathbf{u}^{k,\tau}-\alpha^{u}\mathbf{X}(\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1})]_{\mathbf{\underline{q}}^{k}}^{\mathbf{\overline{q}}^{k}}.$ (14)
The remaining challenge is to compute the gradient in a distributed manner. To achieve this, we exploit the underlying tree topology of radial distribution networks, which enables information propogation in a backward-forward manner. In [35], this procedure is derived using LDU factorization. While backward-forward strategies have been used on tree graphs, our contribution lies in leveraging them to solve the non-Euclidean projection problem, completing a distributed voltage control scheme that optimally coordinates DERs on a fast timescale, operates in a feedback-based manner, and preserves end-users’ privacy by keeping voltage data local. For notational simplicity, we denote $\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1}$ as $\boldsymbol{\omega}^{k,\tau}$ and denote its $i$-th element as $\omega_{i}^{k,\tau}$, which is locally known by node $i$.
In the backward step, for each node $i$, upon collecting $\xi_{j}^{k,\tau}$ for all $j\in\mathcal{C}_{i}$ where $\mathcal{C}_{i}$ represents the set of child nodes of $i$, it sends to its unique parent node
$\xi_{i}^{k,\tau}=\sum_{j\in\mathcal{C}_{i}}\xi_{j}^{k,\tau}+\omega_{i}^{k,\tau}.$ (15)
Note that this step runs recursively from the farthest nodes up to but not including the closest nodes to the point of common coupling (PCC), where $\xi_{j}^{k,\tau}$ is recursively defined. In a nutshell, $\xi_{i}^{k,\tau}$ represents the sum of $\omega_{n}^{k,\tau}$ where $n$ belongs to all of its downstream nodes including itself.
In the forward step, for each node $i$, upon receiving $\zeta_{h}^{k,\tau}$ from its unique parent node, it sends to its child node(s)
$\zeta_{i}^{k,\tau}=\zeta_{h}^{k,\tau}+x_{hi}\xi_{i}^{k,\tau}.$ (16)
Similarly, this step runs recursively from the closest nodes to the PCC up to but not including the farthest nodes, where $\zeta_{h}^{k,\tau}$ is recursively defined. Likewise, $\zeta_{i}^{k,\tau}$ represents the sum of $x_{mn}\xi_{n}^{k,\tau}$ where $n$ belongs to all of its upstream nodes including itself.
Based on the information received from its child node(s) and parent node, each node $i$ can then determines its entry in $\mathbf{X}\boldsymbol{\omega}^{k,\tau}$ as
$[\mathbf{X}\boldsymbol{\omega}^{k,\tau}]_{i}=\zeta_{h}^{k,\tau}+x_{hi}\xi_{i}^{k,\tau}.$ (17)
Note that in this process, each node only communicates with its physical neighbours once. The overall nested feedback optimization algorithm is summarized in Algorithm 1.
Data: $\mathbf{q}^{0}$, $\boldsymbol{\lambda}^{0}$, $\boldsymbol{\mu}^{0}$
Result: $\mathbf{q}^{k}$, $k\geq 0$
1Set k = 0;
@@ 6       Set $\tau=0$;
8            Use the backward-forward algorithm (15)-(17) to compute $\mathbf{X}(\mathbf{u}^{k,\tau}-\mathbf{\dot{q}}^{k+1})$ (choose $\mathbf{u}^{k,0}$ as $\mathbf{q}^{k}$);
9             Calculate next $\mathbf{u}^{\tau+1}$ via (14);
@@ 13      Set $\mathbf{q}^{k+1}=\mathbf{u}^{T}$;
Remark 3.
In principle, the same approach could be applied to (9c) to develop a distributed strategy. However, this would require sharing dual variables. Due to the correlation between voltage and active power, exposing these variables could enable external entities to infer sensitive load information, which is typically considered private. In contrast, our strategy avoids sharing dual variables.
Remark 4.
The proposed method assumes that each DER has access to topological information, such as knowledge of its immediate neighbors. This information is assumed to be obtained during the algorithm’s planning phase. However, we acknowledge that this may not always be feasible. Therefore, exploring approaches that reduce or eliminate this dependency represents an interesting research direction to enhance the practicality and adaptability of the algorithm.
Remark 5.
The addition or removal of DERs may require reconfiguration of the communication graph between agents. However, if each node is interpreted as a virtual DER—potentially representing an aggregation point in the network—the communication structure can remain unchanged. Under this setting, the presence or absence of individual physical DERs does not necessarily impact the agent-level communication.
@@ III-C Convergence analysis
The inner loop represents a standard Euclidean gradient projection algorithm to solve the quadratic program (13). Note that the backward-forward step is only used to gather gradient information; each agent then performs a projected gradient descent step. Consequently, $\mathbf{u}^{\tau}$} is a converging sequence by invoking Proposition 2.3.2 in [29] with $0<\alpha^{u}<2/\lambda_{max}(\mathbf{X})$ where $\lambda_{max}(\mathbf{X})$ is the largest eigenvalue of $\mathbf{X}$. Since we cannot choose $T$ to be arbitrarily large in a practical system, this algorithm will generate an approximate solution to (13). When $T$ is chosen sufficiently large, the algorithm will converge arbitrarily close to the optimum of (13) and the approximation error will be negligible. In Section IV, we demonstrate its convergence with numerical results.
Under the two assumptions, the overall nested algorithm represents a scaled PDGP algorithm to solve (6). By a transformation of variables defined by: $\mathbf{q}^{\prime}=\mathbf{X}^{\frac{1}{2}}\mathbf{q}$ and $\mathbf{v}^{\prime}=\gamma\mathbf{v}$ where $\gamma=\sqrt{\frac{\alpha^{d}}{\alpha}}$ is the squared root of the ratio between the dual and primal step sizes, the problem (6) can be cast as an equivalent quadratic program in the space of $\mathbf{q}^{\prime}$ and $\mathbf{v}^{\prime}$ as in (18).
@@
which are equivalent to (9a)-(9b) and (11)-(13) [29, Eqs. (2.37) and (2.38)]. Invoking Theorem 3 in [2] with $\alpha$ chosen using [2, Eq. (34)], (19) then converges to the optimizer of the regularized saddle-point problem of (18). Consequently, our proposed nested algorithm also converges to the optimizer of the regularized saddle-point problem of (6).
@@
While the linear voltage-reactive power relation is assumed to synthesize the controller and analyze its convergence, AC power flow is run in the simulation to calculate the actual voltages. Such an assumption was also made in the proofs of [3, 27].